OpenStudy (kabase1234):

Please help! I'll fan and medal Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (kabase1234):

an = 7 • (-3)n + 1 an = 7 • 3n - 1 an = 7 • (-3)n - 1 an = 7 • 3n

OpenStudy (kabase1234):

@freckles

OpenStudy (kabase1234):

@phi

OpenStudy (phi):

I think there are formulas to figure this out Do you have notes on this stuff ?

OpenStudy (kabase1234):

No this is new

OpenStudy (kabase1234):

Can you still help? It's my last question

OpenStudy (phi):

you could test each formula if you replace n with 2 you should get -21 do any of the formulas give -21 when n=2 ?

OpenStudy (kabase1234):

C does

OpenStudy (kabase1234):

@phi